solving applications involving variation
simple interest. The simple interest I on an account varies jointly as the time t and the principal P. After one quarter (3 months) an $8000 principal earned $100 in interest. HOw much would a $10000 principal earn in 5 months?
Help, been working on this one problem for over an hour.

Answer by scott8148(6628)

I = k * t * P ___ the constant of proportionality would probably be the monthly interest rate

100 = k * 3 * 8000 ___ k = 1 / 240

I = (1/240) * 5 * 10000